# Point-MoE: Towards Cross-Domain Generalization in 3D Semantic Segmentation via Mixture-of-Experts

### v1

@@ 4.4 Analysis of MoE
Figure 2: Training Loss and Validation mIoU over time. Left: training loss curves for joint training across multiple domains. Right: validation mIoU on indoor datasets. Point-MoE-L achieves competitive performance to PPT, and both significantly outperform the baseline PTv3-L.
Training efficiency and validation mIoU. Fig. 2 shows the training loss and validation mIoU curves for four models: the baseline PTv3-L, its improved variant PTv3-Mix-LN (which incorporates mixed-domain batches and LayerNorm), PPT-L [80], and our proposed Point-MoE-L. All models are trained from scratch across multiple domains. Point-MoE-L converges faster and achieves strong validation mIoU without using explicit dataset labels, matching or exceeding the performance of PPT-L trained with ground-truth domain labels. While all models reach similar training loss, only Point-MoE-L and PPT-L generalize effectively—reinforcing that low training loss is not indicative of strong cross-domain performance. On ScanNet, Structured3D, Matterport3D, and nuScenes, Point-MoE-L shows consistent improvement and avoids early plateaus seen in PPT-L, especially on Structured3D, suggesting stronger long-term learning. PTv3-L fails to generalize and exhibits unstable validation curves. PTv3-Mix-LN improves stability over PTv3-L significantly, but still underperforms Point-MoE-L and PPT-L.
Figure 3: Expert Choices Visualization. Different color represents a different expert. (a) shows a specific expert focusing on edges across the scene, revealing spatially coherent routing. (b) and (c) highlight expert selections in indoor scenes, where semantically related regions (e.g., chair, desk) are consistently routed to the same expert. (d) displays an outdoor scene with sparse LiDAR points, yet the model still demonstrates meaningful routing despite limited geometry.
@@
Expert choice visualization. Fig. 3 showcases expert assignments for a validation scene at selected layers. In (a), we observe that early encoder layers rely heavily on geometric cues for routing. For instance, green experts are consistently activated along object boundaries such as the edges of desks and chairs, while red experts dominate flat surfaces. In (b) and (c), the decoder layers exhibit more semantically meaningful expert selection—likely due to their proximity to the loss function—with distinct experts attending to objects like desks, chairs, floors, and walls. In (d), we examine an outdoor scene with sparse LiDAR data. Despite limited geometric structure, the model still organizes routing meaningfully: nearby points are routed to blue experts, while farther points activate red experts. We include more visualizations in the appendix for completeness. We also note occasional visual artifacts where isolated points are assigned different experts than their neighbors, which may be related to PTv3’s architectural choices such as point serialization or positional encoding.
@@
Figure 4: Expert routing behavior across domains. The left panel shows the most frequently used expert paths through each encoder (E-) and decoder (D-) layer, with feature channel sizes marked in parentheses, illustrating the variation in expert selection for tokens from different datasets. The router tends to assign experts based on the input domain, particularly in the decoder MoE layers. To quantify this, we compute the Jensen-Shannon Divergence (JSD) between the expert selection distributions of different datasets at each MoE layer.
Token pathways. To understand how Point-MoE adapts to diverse domains, we analyze expert routing behavior at the token level. Specifically, we track the top-1 expert assignment for each token across all MoE layers and construct full routing trajectories from the final layer back to the first. We then identify the top 100 most frequent expert paths based on their occurrence across all tokens. As shown in Fig. 4, encoder expert paths are substantially less diverse than those in the decoder, indicating that encoder layers perform more domain-agnostic processing. Interestingly, we observe a sparse routing pattern in deeper encoder layers. This may be attributed not to feature reuse, but to the U-Net-style design where features are spatially deep but token sparsity increases, resulting in reduced variability in routing decisions.
@@
When examining domain-level trends, we find that certain dataset pairs—such as SemanticKITTI and nuScenes or ScanNet and Structured3D—share similar expert pathways, suggesting that Point-MoE implicitly clusters domains with related geometric or semantic structures. To quantify these observations, we compute the Jensen-Shannon Divergence (JSD) [46] between expert selection distributions across datasets at each MoE layer. JSD [46] is an entropy-based measure of divergence between expert routing distributions across datasets, weighted by their token proportions; its formal definition is provided in the supplementary. As shown in Fig. 4 (right), decoder layers exhibit significantly higher JSD, indicating stronger domain-specific specialization. Several encoder layers also display nontrivial JSD, underscoring the benefit of placing MoE throughout the network.
@@ 5 Discussions
This work embraces the “bitter lesson” [66] in AI: scalable generalization emerges from flexible architectures trained on diverse data, rather than hand-engineered domain priors [68]. We introduce Point-MoE, a mixture-of-experts architecture for 3D understanding that dynamically routes tokens to specialized experts across both indoor and outdoor domains. Our experiments demonstrate that Point-MoE not only achieves competitive performance on seen datasets but, crucially, also generalizes effectively to novel domains without explicit domain labels or supervision. Through careful analysis of expert routing behavior, we show that the model self-organizes around both geometric and semantic structures, adapting its computation pathways to the nature of the input. These results highlight the potential of sparse, modular computation for robust and scalable 3D scene understanding. We provide more analysis of Point-MoE and discuss our limitations in the supplementary material.
@@ 6 Acknowledgement
The authors gratefully acknowledge Research Computing at the University of Virginia for providing the computational resources and technical support that made the results in this work possible. Zezhou Cheng acknowledges the Adobe Research Gift, as well as GPU support provided by the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program and the National Artificial Intelligence Research Resource (NAIRR) Pilot program.
@@ A.1 Limitations
Our work is limited by computational resource constraints, which restrict the scale of both the Point-MoE architecture and the training datasets. While our results already demonstrate strong generalization across diverse 3D domains, larger models and a broader training corpus could further enhance performance. Additionally, our expert routing relies solely on the standard gating loss without auxiliary regularization for load balancing or diversity, which may lead to suboptimal expert specialization. Finally, the dynamic routing mechanism incurs nontrivial communication and synchronization overhead compared to dense backbones like Point Transformer V3, highlighting the need for more efficient expert parallelism or conditional computation in future work.
@@ Appendix B Additional Experiments
B.1 Experiment Settings
Data. We compare a diverse set of 3D semantic segmentation datasets spanning both indoor and outdoor environments. ScanNet [21] is a large-scale indoor RGB-D dataset reconstructed from real-world scans of apartments and offices. S3DIS [2] contains real indoor scans of commercial buildings, providing richly annotated scenes with diverse architectural structures. Structured3D [88] is a synthetic dataset based on photo-realistic rendered scenes from CAD models, offering clean labels and dense annotations. Matterport3D [12] consists of real RGB-D scans captured in a wide range of residential environments, covering multiple floors and room types.
@@ B.1 Experiment Settings
Training. Our default joint training strategy and Point-MoE-L configs are summarized in Tab. 8. We experiment with two batching schemes: (1) single batch, where each iteration samples points exclusively from one dataset, and (2) mix batch, where samples are drawn from multiple datasets with equal probabilities. We find that larger batch sizes lead to more stable expert assignment and better convergence behavior in the Mixture-of-Experts architecture. All models are trained from scratch on the joint multi-domain datasets without any fine-tuning stage.
Table 6: Categorical settings of Indoor Datasets.
@@
Table 7: Categorical settings of Outdoor Datasets.
@@
iters 120k iters 160k expert capacity 2H
@@ B.2 Additional Results and Analysis
Domain Detector. To enable domain-aware models at inference time—where ground-truth dataset labels are unavailable—we train a lightweight classifier to predict the domain of an input point cloud. This domain detector serves a purely functional role: it allows domain-conditioned variants of Point-MoE and other baselines to operate in realistic deployment scenarios where users may supply point clouds from unknown or mixed sources. During training, we observe that strong data augmentations (e.g., grid dropout or random cropping) reduce the classifier’s reliability by masking domain-specific cues. To preserve dataset-level characteristics critical for this task, we disable most augmentations and apply only minimal point cropping for efficiency. The domain detector converges quickly, with accuracy stabilizing after just a few epochs. As shown in Fig. 5, it achieves high classification accuracy—99.2% for indoor-only and 95.5% for indoor-outdoor settings—making it suitable for practical use.
@@
Does Point-MoE work on single dataset?
Table 9: Performance of Point-MoE vs. PTv3 on individual datasets. We evaluate both models independently on each dataset to examine whether Point-MoE offers benefits even without cross-domain training. Point-MoE performs comparably to PTv3.
Method ScanNet Structured3D Matterport3D nuScenes Waymo Average
PTv3 75.01 79.37 53.89 77.16 70.20 71.13
Point-MoE 75.63 79.29 53.47 78.24 69.50 71.23
We evaluate Point-MoE without cross-domain supervision by training and testing independently on three indoor datasets (ScanNet [21], Structured3D [88], Matterport3D [12]) and two outdoor datasets (nuScenes [10], Waymo [65]). On ScanNet and nuScenes, Point-MoE outperforms PTv3 by +0.62 and +1.08, respectively. On Structured3D (evaluated on its test split), performance is nearly identical (–0.08). On Matterport3D (test split) and Waymo (validation split), Point-MoE underperforms PTv3 by –0.42 and –0.70, respectively. On average, Point-MoE achieves 71.23 vs. 71.13 for PTv3 (+0.10). These results suggest that expert specialization can emerge in single-domain settings, but the overall improvement is small and not consistent across datasets. We consider the findings inconclusive.
We hypothesize that Point-MoE may require different configurations or training strategies when applied to a single dataset. In particular, the current gating and load balancing mechanisms are optimized for multi-domain settings, where diverse token distributions benefit from expert diversity. In contrast, single-domain training may benefit from fewer experts, a larger top-$k$, or stronger regularization to avoid expert under-utilization.
Table 10: Joint Training Results on Indoor 3D Semantic Segmentation without Precise Evaluator. Each model is trained jointly on three indoor datasets and evaluated under a domain-agnostic setting without relying on ground-truth domain labels during inference. We report performance when models are required to predict across all domains. To evaluate generalization with a single “generic“ label, results marked (G) use this label for both seen and unseen domains. Unmarked results rely on domain detector predictions..
Methods Params Activated Mix Batch Layer. ScanNet Structured3D S3DIS Matterport3D
Val Val Val Val
@@
PTv3-ScanNet [78] 46M 46M 74.4 19.4 8.5 37.9
PTv3-Structured3D [78] 46M 46M 20.0 78.0 7.0 19.5
PTv3-S3DIS [78] 46M 46M 4.9 3.4 66.1 5.2
PTv3-Matterport3D [78] 46M 46M 52.7 23.0 9.5 53.6
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Table 11: Cross-Domain 3D Semantic Segmentation on Indoor and Outdoor Scenes without Precise Evaluator. Models are trained jointly on both indoor and outdoor datasets under domain-agnostic supervision and domain-aware joint training settings. (G): Using the “generic“ label during evaluation (see definition above).
Methods Params Activated Mix Layer. ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo
Val Val Val Val Val Val Val
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
Table 12: Joint Training Performance on Indoor 3D Semantic Segmentation with Ground-Truth Domain Labels. We additionally evaluate models trained on indoor datasets using domain-aware supervision, where explicit ground-truth domain labels are provided during training to guide the learning process.
Methods Params Activated Mix Batch ScanNet Structured3D S3DIS
Val Val Test Val
Domain-aware joint training with G.T. labels
PPT-S [80, 74] 47M 47M 74.7 60.9 65.5 71.6
PPT-L [80, 74] 98M 98M 74.6 60.9 62.6 69.2
Point-MoE-S 59M 52M 76.7 66.7 72.3 71.9
Point-MoE-L 100M 60M 75.6 69.7 71.3 69.6
Point-MoE-G-S 59M 52M 75.7 69.1 70.8 68.1
Point-MoE-G-L 100M 60M 76.7 68.1 69.6 67.1
Table 13: Joint Training Performance on Indoor-Outdoor 3D Semantic Segmentation with Ground-Truth Domain Labels. We evaluate models trained on indoor-outdoor datasets using domain-aware supervision with ground truth labels.
Methods Params Activated Mix ScanNet Structured3D S3DIS nuScenes SemKITTI
Val Val Test Val Val Val
Domain-aware joint training with G.T. labels
PPT-L [80, 74] 98M 98M 76.7 67.2 66.0 71.2 72.6 66.9
Point-MoE-L 100M 60M 76.8 69.6 70.1 70.6 72.4 64.5
Point-MoE-G-L 100M 60M 77.0 74.9 73.2 71.6 73.0 64.3

### v2

@@ 4.4 Analysis of MoE
Dataset clustering emerges in Point-MoE. Fig. 2 shows t-SNE visualizations [70] of encoder and decoder features from PTv3, PPT [82], and Point-MoE. PTv3 exhibits no clear separation in either space, indicating limited capacity to form dataset-specific representations. PPT, by contrast, produces sharply separated clusters at both encoder and decoder levels, reflecting its reliance on dataset labels during training. While this aids seen-datasets performance, such rigid partitioning may hinder generalization capabilities as highlighted in Tab. 2. Point-MoE displays a different pattern: encoder features remain mixed across datasets, supporting shared representation learning, while the decoder disentangles dataset-specific structures. This apparent division of labor suggests that Point-MoE performs implicit dataset inference, without requiring explicit labels. Crucially, on the zero-shot Matterport3D dataset, Point-MoE aligns samples with the most related seen-dataset clusters (e.g., ScanNet), which likely contributes to its strong generalization. Overall, these results show that Point-MoE can self-organize its experts and adapt to novel environments by effectively transferring knowledge across diverse training datasets.
Figure 3: Expert choice visualization. Different colors within each image indicate different experts, showing that Point-MoE self-organizes point routing based on spatial and semantic cues across different layers. (a) shows one expert (green) focusing on edges, suggesting spatially aware mid-level routing in the encoder. (b) and (c) show that semantically related regions (e.g., chairs and desks) are consistently assigned to the same expert, even across datasets. (d) shows meaningful routing in an outdoor scene with sparse LiDAR points despite lower point density and more irregular geometry.
@@
Expert choice visualization. Fig. 3 visualizes expert assignments for validation scenes at selected encoder and decoder layers. In (a), the encoder layer primarily routes points according to local geometric structure. In particular, one expert (green) is consistently activated around object boundaries and thin structural regions, such as the edges of desks and chairs, while other experts dominate flatter surfaces, suggesting spatially aware mid-level routing in the encoder. In (b) and (c), the decoder layers exhibit more semantically coherent expert selection. Semantically related regions, such as chairs, desks, floors, and walls, are more consistently assigned to the same experts, even across different scenes and datasets, likely because decoder features are closer to the supervision signal and thus encode higher-level semantics. In (d), we examine an outdoor scene with sparse LiDAR points. Despite the lower point density and more irregular geometry, the model still produces meaningful routing patterns rather than random assignments. Nearby foreground regions and farther background regions tend to activate different experts, suggesting that the learned routing generalizes beyond dense indoor scans to more challenging outdoor environments. Additional visualizations are provided in Sec. C.3. We also observe a few isolated points assigned to different experts than their neighbors, possibly due to PTv3 design choices such as point serialization and positional encoding.
Table 5: Inference efficiency. FLOPs per step are measured with a profiler, and peak VRAM reports allocated memory. Relative deltas are computed with respect to PPT-L.
Model FLOPs/step (GFLOPs) Peak VRAM (GiB) $\Delta$ FLOPs vs. PPT-L $\Delta$ VRAM vs. PPT-L
PPT-L 384.4 41.1 0% 0%
PTv3 375.8 37.4  $\downarrow\,2.2\%$  $\downarrow\,9.0\%$
Point-MoE-L 265.7 33.3  $\downarrow\,30.9\%$  $\downarrow\,19.0\%$
@@
Figure 4: Expert routing across datasets. Left: the most frequent expert paths through each encoder (E) and decoder (D) layer, with channel sizes in parentheses, showing that expert selection varies across datasets. The effect is especially pronounced in decoder MoE layers. Right: Jensen Shannon Divergence (JSD) between dataset specific expert selection distributions at each MoE layer.
@@
Token pathways. To understand how Point-MoE adapts to diverse datasets, we analyze expert routing at the token level. We track each token’s top-1 expert across all MoE layers and reconstruct full routing trajectories from the final layer back to the first. We then rank expert paths by frequency and focus on the top 100 (Fig. 4). Encoder paths are notably less diverse than decoder paths, suggesting that encoders perform more dataset-agnostic processing. We also find sparse routing in deeper encoder layers, likely due not to feature reuse but to the U-Net design: features grow deeper spatially while tokens become sparser, reducing routing variability. When examining dataset-level trends, we find that certain dataset pairs, such as SemanticKITTI and nuScenes or ScanNet and Structured3D, share similar expert pathways, suggesting that Point-MoE implicitly clusters datasets with related geometric or semantic structures. To quantify these observations, we compute the Jensen-Shannon Divergence (JSD) [45] between expert selection distributions across datasets at each MoE layer. JSD [45] is an entropy-based measure of divergence between expert routing distributions across datasets, weighted by their token proportions; its formal definition is provided in the supplementary. As shown in Fig. 4 (right), decoder layers exhibit significantly higher JSD, indicating stronger dataset-specific specialization. Several encoder layers also display nontrivial JSD, underscoring the benefit of placing MoE throughout the network (see Sec. C.2 for more analysis).
Efficiency analysis. As shown in Tab. 5, Point-MoE-L requires only 265.7 GFLOPs per step and 33.3 GiB peak VRAM, reducing compute by $30.9\%$ and memory by $19.0\%$ relative to PPT-L (384.4 GFLOPs and 41.1 GiB). It is also more efficient than PTv3 (375.8 GFLOPs and 37.4 GiB). Despite this lower cost, Point-MoE-L achieves the best average mIoU in both settings, reaching $71.5$ in indoor only joint training and $70.2$ in indoor and outdoor joint training (Tab. 1).
@@ 5 Discussions
This work embraces the “bitter lesson” [68] in AI: scalable generalization emerges from flexible architectures trained on diverse data, rather than hand-engineered domain priors. We introduce Point-MoE, a mixture-of-experts architecture for 3D understanding that dynamically routes tokens to specialized experts across both indoor and outdoor domains. Our experiments demonstrate that Point-MoE not only achieves superior performance on seen datasets but also generalizes effectively to zero-shot datasets without explicit domain labels. Through careful analysis of expert routing behavior, we show that the model self-organizes around both geometric and semantic structures, adapting its computation pathways to the nature of the input. These results highlight the potential of sparse, modular computation for robust and scalable 3D scene understanding. We provide more analysis of Point-MoE and discuss our limitations in the supplementary materials.
@@ 6 Acknowledgement
The authors acknowledge the MathWorks Research Award, Adobe Research Gift, the University of Virginia Research Computing and Data Analytics Center, Advanced Micro Devices AI and HPC Cluster Program, Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, and National Artificial Intelligence Research Resource (NAIRR) Pilot for computational resources, including the Anvil supercomputer (National Science Foundation award OAC 2005632) at Purdue University and the Delta and DeltaAI advanced computing resources (National Science Foundation award OAC 2005572).
@@ A.1 Limitations
Within our compute budget, we present a comprehensive study of MoE for point clouds—covering placement, expansion, routing choices (top-$k$, capacity factor, load balancing), and mixed-domain batching—with consistent gains. However, compute remains the primary bottleneck, capping Point-MoE’s scale (experts, width, depth), training schedule, and number of jointly trained datasets. We therefore leave systematic scaling (e.g., expert count/width, per-stage sharing, alternative routers) to future work. Finally, because 3D semantic segmentation depends on costly human labels, we plan to pursue self-supervised pretraining on diverse unlabeled point clouds to reduce annotation effort and improve generalization. We hope these results serve as a simple, strong baseline and a useful reference point for future works.
@@ Appendix B Additional Experiments
B.1 Experiment Datasets
We compare a diverse set of 3D semantic segmentation datasets spanning both indoor and outdoor environments. ScanNet [19] is a large-scale indoor RGB-D dataset reconstructed from real-world scans of apartments and offices. S3DIS [1] contains real indoor scans of commercial buildings, providing richly annotated scenes with diverse architectural structures. Structured3D [95] is a synthetic dataset based on photo-realistic rendered scenes from CAD models, offering clean labels and dense annotations. Matterport3D [10] consists of real RGB-D scans captured in a wide range of residential environments, covering multiple floors and room types.
@@ B.1 Experiment Datasets
Table 8: Categories of Indoor Datasets.
@@
Table 9: Categories of Outdoor Datasets.
@@
iters 140k iters 180k expert capacity 2H
@@ B.2 Additional Results and Analysis
Domain Detector.
To enable domain-aware models at inference time—where ground-truth dataset labels are unavailable—we train a lightweight classifier to predict the domain of an input point cloud. This domain detector serves a purely functional role: it allows domain-conditioned variants of Point-MoE and other baselines to operate in realistic deployment scenarios where users may supply point clouds from unknown or mixed sources. During training, we observe that strong data augmentations (e.g., grid dropout or random cropping) reduce the classifier’s reliability by masking domain-specific cues. To preserve dataset-level characteristics critical for this task, we disable most augmentations and apply only minimal point cropping for efficiency. The domain detector converges quickly, with accuracy stabilizing after just a few epochs. As shown in Fig. 5, it achieves high classification accuracy—99.2% for indoor-only and 95.5% for indoor-outdoor settings—making it suitable for practical use.
@@
Does Mixture-of-Experts (MoE) work with a sparse convolutional backbone?
Table 11: MoE on a sparse convolutional backbone (OA-CNN). Numbers are mIoU. Rows (1)–(4) use the public OA-CNN implementation (with a language-guided head for cross-dataset evaluation). (1)–(3) train on a single dataset; (4) joint-trains on ScanNet, Structured3D, and S3DIS; (5) replace SparseCNN with MoE-CNN using top-2 routing.
Model ScanNet Structured3D S3DIS Matterport3D Average
(1) OA-CNN-S3DIS 5.70 3.8 65.3 5.4 20.6
(2) OA-CNN-ScanNet 72.6 21.6 8.4 37.4 35.0
(3) OA-CNN-Structured3D 20.8 65.6 6.8 18.9 28.0
(4) OA-CNN 71.6 59.0 60.6 37.8 57.3
(5) OA-CNN-MoE 72.0 65.3 63.0 40.2 60.1
Yes. We integrate MoE into OA-CNN [52] by replacing the SparseCNN layer in the BasicBlock with a mixture of sparse-convolution experts (MoE-CNN), evaluating with top-2 routing. As summarized in Table 11, MoE improves over the joint-trained OA-CNN baseline on seen domains (ScanNet, Structured3D, S3DIS) and the unseen Matterport3D, raising the average from 57.3 to 60.1. This demonstrates that MoE is compatible with sparse convolutional backbones and delivers measurable gains, which further demonstrates MoE’s applicability to architecture beyond transformers.
Table 12: Main Results without Precise Evaluator. Each model is trained jointly on three indoor datasets and evaluated without precise evaluator. * indicates that each training batch combines samples from multiple datasets, methods with ** use mix dataset batch and layernorm [3].
Methods Params Activated ScanNet Structured3D S3DIS Matterport3D Average
Val Val Area5 Val
@@
PTv3-ScanNet [81] 46M 46M 74.4 19.4 8.5 37.9 35.1
PTv3-Structured3D [81] 46M 46M 20.0 78.0 7.0 19.5 31.1
PTv3-S3DIS [81] 46M 46M 4.9 3.4 66.1 5.2 19.9
PTv3-Matterport3D [81] 46M 46M 52.7 23.0 9.5 53.6 34.7
Multi-dataset Joint Training
PTv3-[81] 46M 46M 48.4 57.8 44.1 27.8 44.5
PTv3-S* [81] [81] 46M 46M 67.7 45.9 62.5 34.9 52.8
PTv3-S** [81] 46M 46M 70.0 56.0 66.9 35.6 57.1
PTv3-[81] 97M 97M 57.1 58.1 40.3 31.0 46.6
PTv3-L* [81] 97M 97M 73.3 57.5 67.2 38.4 59.1
PTv3-L** [81] [81] 97M 97M 74.0 67.4 67.8 40.6 62.5
PPT-[82, 75] 47M 47M 74.2 58.2 61.0 24.3 54.4
PPT-[82, 75] 98M 98M 74.7 64.9 62.0 27.8 57.4
Point-MoE-S 59M 52M 74.5 66.9 68.6 40.1 62.5
Point-MoE-L 100M 60M 75.1 69.3 68.5 41.3 63.6
Table 13: Main Results on Indoor and Outdoor Scenes without Precise Evaluator. Models are trained jointly on both indoor and outdoor datasets. indicates that each training batch combines samples from multiple datasets, methods with ∗∗ use mix dataset batch and layernorm [3].
Methods Params Activated ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo Average
Val Val Area5 Val Val Val Val
PTv3-L [81] 97M 97M 55.9 47.6 19.3 42.9 35.3 31.8 12.4 35.0
PTv3-L* [81] 97M 97M 73.4 61.9 61.4 64.7 59.3 38.9 22.8 54.6
PTv3-L** [81] 97M 97M 73.6 61.3 65.1 64.5 59.7 37.1 21.9 54.7
PPT-L [82, 75] 98M 98M 75.2 67.5 61.0 57.2 64.0 20.8 16.0 51.7
Point-MoE-L 100M 60M 76.2 70.1 67.9 68.9 63.3 40.7 23.4 58.6